### v1

@@ II.2.5 DUNE
As recommended in the 2023 P5 report [P5Report], the DUNE Collaboration seeks to construct two 10-kton fiducial liquid argon time projection chambers (LArTPC), deep underground in the Homestake mine in South Dakota [2016arXiv160102984A]. A third LArTPC will follow in Phase II, along with other upgrades. The newly developed Sanford Underground Research Facility (SURF) offers 4800-m water equivalent overburden, and will form the far site for a long baseline neutrino program utilizing a high-energy neutrino beam directed from Fermilab to SURF. The DUNE LArTPC detectors may be sensitive to measurements of the high-energy solar neutrinos via CC interactions on argon, which offers good precision on the spectral shape and, thus, the potential to measure day/night effects [2019PhRvL.123m1803C]. The solar neutrino program at DUNE is limited primarily by backgrounds (the detector is optimized for GeV-scale physics) and energy threshold. Multiple technologies are under consideration for a fourth detector module, with the motivation of expanding the physics program, including improved sensitivity to solar neutrinos. Use of underground argon, or alternative technologies such as organic or water-based scintillators, could preserve the long baseline neutrino sensitivity while also opening up a rich program of low-energy physics.
@@ II.2.6 Future prospects
As experiments grow larger, and capabilities increase, there are also opportunities to leverage detectors designed for other purposes. Noble liquid detectors built primarily for dark matter, or even long baseline neutrino experiments, may offer sensitivity to solar neutrinos. Noble liquids have high light yields, and are transparent to their own scintillation. Coupled with the high voltage under which such detectors are often operated, the combined ionisation and scintillation signatures can be powerful for discriminating signal from background. The low threshold of these detectors may allow detection of neutrinos from the high-energy tail of the 8B spectrum via coherent elastic scattering off a target nucleus. Solid state detectors built purposely for observing this “CENS” signature have been used to detect neutrinos from stopped pion decay [doi:10.1126/science.aao0990].
III The reaction ()
The cross section for the initial reaction in the pp chain (Figure 1), p+p, is too small to be measured in the laboratory. It must be calculated from the standard theory of weak interactions.
@@ III.1 Introduction and terminology
Near the Gamow peak energy 6 keV for temperatures characteristic of the Sun center, the first and second derivatives of the astrophysical factor at zero energy, and , generate 7% and 0.5% corrections, respectively, in Taylor’s series expansion of around . Higher derivative terms are neglected in this review since they only contribute at the level. The recommended values for and will be discussed in Section III.5. Here we focus on .
@@ III.2 Adopted parameters for this review
For the phase-space factor , we use the same value as in SF-II. It comes from the value without radiative corrections, [1969ApJ...155..501B], increased by 1.62% to take into account radiative corrections to the cross section [2003PhRvC..67c5502K]. The main source of uncertainty in arises from neglected diagrams in which the lepton exchanges a weak boson and a photon with different nucleons. These diagrams are estimated to modify by , based on scaling the similar nucleus-dependent correction in superallowed decay [2003PhRvC..67c5502K]. Direct computations of these diagrams were recommended in Solar Fusion II. Here we urge again that this computation be carried out.
@@
For , we use the PDG [2022PTEP.2022h3C01W] value, , whose central value is larger than the 2008 PDG value, , used in SF-II by (or 1.9 ). Naively, this would lead to a increase in the central value of according to Eq. (5). Its effect will be discussed below. For , we take s from the most updated comprehensive analysis of experimental rates with the radiative and Coulomb effects corrected [2020PhRvC.102d5501H]. This value is consistent with s [2009PhRvC..79e5502H] used in SF-II with a larger error.
@@
The dominant uncertainty in comes from the normalized Gamow-Teller (GT) matrix element . Reducing this uncertainty has been the main focus of theoretical work since SF-I. In SF-I, was written as , i.e. the sum of the one- and two-body current matrix elements, and , respectively, with their uncertainties estimated independently. In SF-II, two major steps had contributed to reducing the uncertainty on . The first was a much deeper understanding of the correlation between the uncertainties in and : the overall uncertainty in could be described by a universal parameter that could be fixed by a single measurement. The study of 1998PhRvC..58.1263S demonstrated this phenomenologically in the context of potential-model approaches, while later analysis via effective field theory (EFT) provided a more formal justification [2001PhRvC..63c5501B, 2003PhRvC..67e5206P]. The second step was the use of the precisely known tritium decay rate to fix this universal parameter, as first proposed by 1991PhRvC..44..619C. This has been done in both potential models [1998PhRvC..58.1263S] and in the hybrid EFT approach [2003PhRvC..67e5206P] as explained below.
@@
In SF-II, was determined with three approaches. The first one was the potential model approach. In the most elaborate calculation for the pp fusion process, a comparison of the results for five representative modern potentials designed to accurately reproduce nucleon-nucleon scattering data was carried out [1998PhRvC..58.1263S]. After adjusting the unknown strength of the two-body exchange currents to reproduce , the variation in that otherwise would come from the choice of the phenomenological potential was largely removed. Predictions for five representative high-precision phenomenological potentials fell in a narrow interval . There were additional uncertainties in the three-body potentials and three-body currents in , of the order of , and a 0.5% uncertainty due to effective range parameters for nucleon-nucleon scattering. Hence, the recommended value from the potential model approach was .
@@
Based on the consistent results of the above three approaches, SF-II recommended . The SF-I value from 1969ApJ...155..501B was not re-computed in SF-II. However, with reaching a accuracy, new calculations of and , together with the full pionful chiral EFT (EFT) computations to remove the unknown systematics of the hybrid EFT, were called for in SF-II. These two challenges have been met in SF-III and will be described below.
@@ III.3 Experimental progress on muon capture of the deuteron
In SF-II it was also recommended to carry out the experimental determination of the muon capture rate on deuteron, as proposed in the MuSun experiment [2003nucl.ex...4019K, 2017APS..DNP.CG007S]. This quantity could be used to constrain without the need to rely on the three-body calculation of . The theoretical calculations for the muon capture rate on deuteron have been carried out in EFT [2005PhRvC..72f1001C] (it is possible to impose neutron energy cut to isolate the low-energy neutron events so EFT is applicable [2003nucl.ex...4019K]), chiral hybrid EFT* [2002PhLB..533...25A], phenomenological potential model [2011PhRvC..83a4002M], and recently in EFT (2012PhRvL.108e2502M, 2018PhRvL.121d9901M; 2018PhRvC..98f5506A; 2023FrP....1049919C; 2023PhRvC.107f5502B; PhysRevC.109.035502). However, the MuSun result is yet to be released. On the other hand, we should notice that the muon capture processes happen at a rather large momentum transfer compared to pp fusion. The momentum transfer dependence of the single nucleon axial coupling constant , with , has been recently studied by 2018RPPh...81i6301H, who have provided an experimental determination for the axial charge radius, given by fm2, where is defined by the relation for small . The % uncertainty on has an impact on the ability of the MuSun experiment alone to directly constrain [2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502].
@@ III.4 Progress in calculations since SF II
Below we summarize the calculations performed after SF-II using different approaches.
@@ III.4.1 EFT
In the most recent and comprehensive EFT study, 2023arXiv230403327A compared the above calculations of 2019PhRvL.123a9901M and 2016PhLB..760..584A in detail. In addition to the increase of from using the correct - relation, updating the input parameters to their most recent values increased by in both calculations, mainly due to the increase of from its SF-II value mentioned above. Furthermore, 2019PhRvL.123a9901M received a increase from removing the truncation error of the basis functions which effectively cut off the long distance part of the wave functions, as 2017PhRvC..95c1301A advocated. After these corrections, the difference between 2019PhRvL.123a9901M and 2016PhLB..760..584A was reconciled and the combined result was found to be [2023arXiv230403327A]
@@
Finally, we would like to mention the work of 2022PhRvC.106e5501L, where the power counting of the EFT weak current operator involved in the pp reaction is revisited using renormalization group (RG) invariance as the guideline. In particular, it is argued that the contact two-body axial current proportional to the LEC must appear one order lower than assessed by naive dimensional analysis. Then it can be shown that RG invariance is fulfilled at N2LO. However, the estimate for obtained by 2022PhRvC.106e5501L does not use a value for obtained by fitting , but extracted in order to match the value of obtained in SF-II. Therefore, we will not consider the work of 2022PhRvC.106e5501L in the present evaluation.
@@ III.4.4 Final Recommendation of
The above discussions show that determinations of from EFT in Eq.(6), EFT in Eq.(7), and lattice QCD (although with unquantified systematics) are all consistent with each other. Furthermore, these values are also consistent with the recommended value of SF-II, provided the central value is increased by 444To know the precise shift requires an explicit calculation. However, if the shift is within the range of 0.8-1.0 , the averaged value and remain the same within the significant digits. to to account for the update. Averaging this value with the EFT value in Eq.(6) and the EFT value in Eq.(7) yields with per degree of freedom to be . This shows that the EFT, EFT and SF-II estimates of are all mutually consistent.
In addition, we would like to advocate adding an additional correlated error to account for any input that would tend to move all results in a coordinated way. For example, from SF-II to SF-III, we experienced the shift due to the update of the PDG value of . It is not inconceivable that or other input parameters or physics could change again by similar amounts: the large PDG inflation factor of 2.7 reflects the tension that continues to exist among measurements [2022PTEP.2022h3C01W]. Therefore, assigning an additional correlated error seems reasonable. This is also in line with the subtleties discussed in Secs.III.4.1 and III.4.2 which call for an enlarged error. Therefore, our final recommended value for is
@@ IV.1 Introduction
The 2H(,)3He reaction is the second step in the pp chain (Figure 1). Compared with the reactions mediated by the weak interaction, this reaction occurs much more rapidly. As a result, deuterium is effectively and instantaneously converted to 3He without leaving any observable signature. Therefore, uncertainties in the rate of this reaction do not affect solar energy generation.
@@
Another astrophysical scenario where the 2H(,)3He reaction plays a key role is Big Bang Nucleosynthesis (BBN), responsible for the production of light elements during the first few minutes of the Universe. Among these elements, deuterium is an excellent indicator of cosmological parameters because its primordial abundance is the most sensitive to the baryon density and critically depends on the radiation density of the early Universe. The reactions involved in the synthesis of deuterium are: production via the well known )2H process and destruction via the 2H(2H,)3He, 2H(2H,)3H and 2H(,)3He reactions [2020JCAP...03..010F].
Since the comprehensive review performed in SF II, there have been both new measurements as well as advances in the theoretical and phenomenological analysis of the 2H(,)3He reaction. The new experimental results have been determined with accelerator-based measurements of the cross section [2019EPJA...55..137T, 2020Natur.587..210M, 2021PhRvC.103d5805T], as well as plasma-based, intertial confinement fusion measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M]. On the theoretical side, there have been advances in ab initio calculations, where nucleons are the fundamental degrees of freedom interacting among themselves and with the external electromagnetic probe. Finally, Bayesian analysis methods have been used to model the energy dependence of the factor, starting from the ab initio predictions and applying a polynomial approximation to it.
We review these updates and provide recommended values and uncertainties of the factor over the energy range of interest for Solar fusion, based upon a Bayesian averaging of various models.
@@ IV.2 Data sets used for the present review
The 2H(,)3He reaction has a Q value of 5.5 MeV and proceeds through the direct capture mechanism. Different experimental approaches were followed to measure its cross section. Tisma et al. [2019EPJA...55..137T] irradiated deuterated titanium targets with a proton beam and detected the -rays with two high-purity germanium (HPGe) detectors placed at different angles. The final factor is provided at four energies in the 47 - 210 keV range, with approximately 15% uncertainty. More recently, the LUNA Collaboration performed a measurement underground in the Gran Sasso Laboratories, exploiting the six orders of magnitude suppression of the cosmic radiation background [2021FrASS...7..119Z, 2018IJMPA..3343010C]. A windowless deuterium gas target was used and the -rays emitted by the 2H(,)3He reaction were detected by a large HPGe detector at 90∘ with respect to the beam axis. Great care was taken to minimize all sources of systematic uncertainties in the factor data, resulting in high precision and accuracy at the 3% level [2020EPJA...56..144M, 2020Natur.587..210M]. These new results provided stringent constraints on cosmological parameters obtained by comparing the precise primordial deuterium abundance predictions of the standard BBN model with astronomical observations [2018ApJ...855..102C]. A deeper discussion of the LUNA results and their implications can be found in [2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y, 2021ApJ...923...49M]. Presently a study of the 2H(,)3He angular distribution is ongoing. By studying the shape of the peak produced in the -spectra by the reaction, it is possible to obtain the angular distribution; a dedicated paper containing the results will be published soon. Finally, a new measurement was performed at the Helmholtz-Zentrum Dresden-Rossendorf in the 300-1000 keV energy range [2021PhRvC.103d5805T] using implanted deuterium targets on tantalum backings and two HPGe detectors. The results are affected by very large systematic uncertainties and they show a different trend of the factor at energies above 300 keV with respect to the fit of the LUNA data of 2020Natur.587..210M. To investigate the tension between the LUNA fit and the more recent high-energy data of 2021PhRvC.103d5805T, a new measurement of the 2H(,)3He reaction is planned at the Felsenkeller laboratory in Germany [2019sone.conf..249B].
@@
In addition to the new accelerator-based results, two recent sets of measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] have also been performed using the inertial confinement fusion plasma-based platform [2017PhPl...24d1407G, 2022FrP....10.4339M, 2023FrP....1180821G], which has recently begun to be exploited for this type of work, with initial results on other reactions presented in, e.g., [2016PhRvL.117c5002Z, 2017NatPh..13.1227C]. Both of these measurements, which were done at the OMEGA laser facility (see section LABEL:subsec:WG8:Plasma), used laser-driven implosions of spherical plastic-shell capsules filled with H2D2 gas, and measured the emitted -rays using a gas-Cherenkov-detector [2014RScI...85kE124H] that was calibrated applying the technique described in 2019RScI...90l3504Z. The initial experiment [2020PhRvC.101d2802Z] obtained good statistics at an energy of 16 keV by making several repeated measurements, with a final statistical uncertainty of 6% and a systematic one of 17% (dominated by uncertainty in the absolute calibration of the detector). The second experiment [2022FrP....10.4339M] obtained data at three different energies in the 17-37 keV region, with comparable systematic uncertainty, but with larger statistical uncertainty due to fewer repeated measurements. The results obtained on this unique platform agree within error bars with the accelerator-based measurements.

### v2

@@ II.2.5 DUNE
As recommended in the 2023 P5 report [P5Report], the DUNE Collaboration seeks to construct two 10-kton fiducial liquid argon time projection chambers (LArTPC), deep underground in the Homestake mine in South Dakota [2016arXiv160102984A]. A third LArTPC will follow in Phase II, along with other upgrades. The newly developed Sanford Underground Research Facility (SURF) offers 4800-m water equivalent overburden, and will form the far site for a long baseline neutrino program utilizing a high-energy neutrino beam directed from Fermilab to SURF. The DUNE LArTPC detectors may be sensitive to measurements of the high-energy solar neutrinos via charged current interactions on argon, which offers good precision on the spectral shape and, thus, the potential to measure day/night effects [2019PhRvL.123m1803C]. The solar neutrino program at DUNE is limited primarily by backgrounds (the detector is optimized for GeV-scale physics) and energy threshold. Multiple technologies are under consideration for a fourth detector module, with the motivation of expanding the physics program, including improved sensitivity to solar neutrinos. Use of underground argon, or alternative technologies such as organic or water-based scintillators, could preserve the long baseline neutrino sensitivity while also opening up a rich program of low-energy physics.
@@ II.2.6 Future prospects
As experiments grow larger, and capabilities increase, there are also opportunities to apply detectors designed for other purposes to the detection of solar neutrinos. Noble liquid detectors built primarily for dark matter searches and long-baseline neutrino experiments have now reached the size and background levels to be sensitive to solar neutrinos [2024arXiv240802877A]. The detection of neutrinos from the high-energy tail of the 8B spectrum via coherent elastic neutrino-nucleus scattering (“CENS”) is particularly interesting because it is a purely neutral-current process sensitive to the total flux of active neutrinos (above the experimental threshold, about 8 MeV) from the Sun. Only the SNO experiment has had such a capability. Initial results reported in [2024arXiv240802877A] are in agreement, with limited statistical precision. Solid state detectors built purposely for observing CENS have observed neutrinos from stopped pion decay at the theoretically expected rate [doi:10.1126/science.aao0990].
The 1HH reaction ()
The cross section for the initial reaction in the pp chain (Figure 1), , is too small to be measured in the laboratory. It must be calculated from the standard theory of weak interactions.
@@ III.1 Introduction and terminology
Near the Gamow peak energy 6 keV for temperatures characteristic of the solar core, the first and second derivatives of the astrophysical factor at zero energy, and , generate 7% and 0.5% corrections, respectively, in Taylor’s series expansion of around . Higher derivative terms are neglected in this review since they only contribute at the level. The recommended values for and will be discussed in Section III.5. Here we focus on .
@@ III.2 Adopted parameters for this review
For the phase-space factor , we use the same value as in 2011RvMP...83..195A. It comes from the value without radiative corrections, [1969ApJ...155..501B], increased by 1.62% to take into account radiative corrections to the cross section [2003PhRvC..67c5502K]. The main source of uncertainty in arises from neglected diagrams in which the lepton exchanges a weak boson and a photon with different nucleons. These diagrams are estimated to modify by , based on scaling the similar nucleus-dependent correction in superallowed decay [2003PhRvC..67c5502K]. Direct computations of these diagrams were recommended in Solar Fusion II. Here we urge again that this computation be carried out.
@@
For , we use the value from the Particle Data Group (PDG) compilation [2022PTEP.2022h3C01W], , whose central value is larger than the 2008 PDG value, , used in 2011RvMP...83..195A by (or 1.9 ). Naively, this would lead to a increase in the central value of according to Eq. (5). Its effect will be discussed below. For , we take s from the most updated comprehensive analysis of experimental rates with the radiative and Coulomb effects corrected [2020PhRvC.102d5501H]. This value is consistent with s [2009PhRvC..79e5502H] used in 2011RvMP...83..195A with a larger error.
@@
The dominant uncertainty in comes from the normalized Gamow-Teller (GT) matrix element . Reducing this uncertainty has been the main focus of theoretical work since 1998RvMP...70.1265A. In 1998RvMP...70.1265A, was written as , i.e. the sum of the one- and two-body current matrix elements, and , respectively, with their uncertainties estimated independently. In 2011RvMP...83..195A, two major steps had contributed to reducing the uncertainty on . The first was a much deeper understanding of the correlation between the uncertainties in and : the overall uncertainty in could be described by a universal parameter that could be fixed by a single measurement. The study of 1998PhRvC..58.1263S demonstrated this phenomenologically in the context of potential-model approaches, while later analysis via effective field theory (EFT) provided a more formal justification [2001PhRvC..63c5501B, 2003PhRvC..67e5206P]. The second step was the use of the precisely known tritium decay rate to fix this universal parameter, as first proposed by 1991PhRvC..44..619C. This has been done in both potential models [1998PhRvC..58.1263S] and in the hybrid EFT approach [2003PhRvC..67e5206P] as explained below.
@@
In 2011RvMP...83..195A, was determined with three approaches. The first one was the potential model approach. In the most elaborate calculation for the pp fusion process, a comparison of the results for five representative modern potentials designed to accurately reproduce nucleon-nucleon scattering data was carried out [1998PhRvC..58.1263S]. After adjusting the unknown strength of the two-body exchange currents to reproduce , the variation in that otherwise would come from the choice of the phenomenological potential was largely removed. Predictions for five representative high-precision phenomenological potentials fell in a narrow interval . There were additional uncertainties in the three-body potentials and three-body currents in , of the order of , and a 0.5% uncertainty due to effective range parameters for nucleon-nucleon scattering. Hence, the recommended value from the potential model approach was .
@@
Based on the consistent results of the above three approaches, 2011RvMP...83..195A recommended . The 1998RvMP...70.1265A value from 1969ApJ...155..501B was not re-computed in 2011RvMP...83..195A. However, with reaching a accuracy, new calculations of and , together with the full pionful chiral EFT (EFT) computations to remove the unknown systematics of the hybrid EFT, were called for in 2011RvMP...83..195A. These two challenges have been met in SF III and will be described below.
@@ III.3 Experimental progress on muon capture of the deuteron
In 2011RvMP...83..195A it was also recommended to carry out the experimental determination of the muon capture rate on deuteron, as proposed in the MuSun experiment [2003nucl.ex...4019K, 2017APS..DNP.CG007S]. This quantity could be used to constrain without the need to rely on the three-body calculation of . The theoretical calculations for the muon capture rate on deuteron have been carried out in EFT [2005PhRvC..72f1001C] (it is possible to impose neutron energy cut to isolate the low-energy neutron events so EFT is applicable [2003nucl.ex...4019K]), chiral hybrid EFT* [2002PhLB..533...25A], phenomenological potential model [2011PhRvC..83a4002M], and recently in EFT (2012PhRvL.108e2502M, 2018PhRvL.121d9901M; 2018PhRvC..98f5506A; 2023FrP....1049919C; 2023PhRvC.107f5502B; PhysRevC.109.035502). However, the MuSun result is yet to be released. On the other hand, we should notice that the muon capture processes happen at a rather large momentum transfer compared to pp fusion. The momentum transfer dependence of the single nucleon axial coupling constant , with , has been recently studied by 2018RPPh...81i6301H, who have provided an experimental determination for the axial charge radius, given by fm2, where is defined by the relation for small . The % uncertainty on has an impact on the ability of the MuSun experiment alone to directly constrain [2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502].
@@ III.4 Progress in calculations since SF II
Below we summarize the calculations performed after 2011RvMP...83..195A using different approaches.
@@ III.4.1 EFT
In the most recent and comprehensive EFT study, 2023arXiv230403327A compared the above calculations of 2019PhRvL.123a9901M and 2016PhLB..760..584A in detail. In addition to the increase of from using the correct - relation, updating the input parameters to their most recent values increased by in both calculations, mainly due to the increase of from its 2011RvMP...83..195A value mentioned above. Furthermore, 2019PhRvL.123a9901M received a increase from removing the truncation error of the basis functions which effectively cut off the long distance part of the wave functions, as 2017PhRvC..95c1301A advocated. After these corrections, the difference between 2019PhRvL.123a9901M and 2016PhLB..760..584A was reconciled and the combined result was found to be [2023arXiv230403327A]
@@
Finally, we would like to mention the work of 2022PhRvC.106e5501L, where the power counting of the EFT weak current operator involved in the pp reaction is revisited using renormalization group (RG) invariance as the guideline. In particular, it is argued that the contact two-body axial current proportional to the LEC must appear one order lower than assessed by naive dimensional analysis. Then it can be shown that RG invariance is fulfilled at N2LO. However, the estimate for obtained by 2022PhRvC.106e5501L does not use a value for obtained by fitting , but extracted in order to match the value of obtained in 2011RvMP...83..195A. Therefore, we will not consider the work of 2022PhRvC.106e5501L in the present evaluation.
@@ III.4.4 Final Recommendation of
The above discussions show that determinations of from EFT in Eq.(6), EFT in Eq.(7), and lattice QCD (although with unquantified systematics) are all consistent with each other. Furthermore, these values are also consistent with the recommended value of 2011RvMP...83..195A, provided the central value is increased by 555To know the precise shift requires an explicit calculation. However, if the shift is within the range of 0.8-1.0 , the averaged value and remain the same within the significant digits. to to account for the update. Averaging this value with the EFT value in Eq.(6) and the EFT value in Eq.(7) yields with per degree of freedom to be . This shows that the EFT, EFT and 2011RvMP...83..195A estimates of are all mutually consistent.
In addition, we would like to advocate adding an additional correlated error to account for any input that would tend to move all results in a coordinated way. For example, from 2011RvMP...83..195A to SF III, we experienced the shift due to the update of the PDG value of . It is not inconceivable that or other input parameters or physics could change again by similar amounts: the large PDG inflation factor of 2.7 reflects the tension that continues to exist among measurements [2022PTEP.2022h3C01W]. Therefore, assigning an additional correlated error seems reasonable. This is also in line with the subtleties discussed in Secs.III.4.1 and III.4.2 which call for an enlarged error. Therefore, our final recommended value for is
@@ IV.1 Introduction
The 2H(,)3He reaction is the second step in the pp chain (Figure 1). Compared with the reactions mediated by the weak interaction, this reaction occurs much more rapidly. Consequently, on the time scales relevant to solar energy generation, deuterium is effectively converted to 3He instantaneously and thus it is only sensitive to the value of the reaction and not the uncertainties of the rate.
@@
Another astrophysical scenario where the 2H(,)3He reaction plays a key role is Big Bang Nucleosynthesis (BBN), responsible for the production of light elements during the first few minutes of the Universe. Among these elements, deuterium is an excellent indicator of cosmological parameters because its primordial abundance is the most sensitive to the baryon density and critically depends on the radiation density of the early Universe, see for example a review of BBN [2016RvMP...88a5004C]. The reactions involved in the synthesis of deuterium are: production via the well known )2H process and destruction via the 2H(2H,)3He, 2H(2H,)3H and 2H(,)3He reactions [2020JCAP...03..010F].
Since the comprehensive review performed in 2011RvMP...83..195A, there have been both new measurements as well as advances in the theoretical and phenomenological analysis of the 2H(,)3He reaction. The new experimental results have been determined with accelerator-based measurements of the cross section [2019EPJA...55..137T, 2020Natur.587..210M, 2021PhRvC.103d5805T], as well as plasma-based, intertial confinement fusion measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] and are reviewed in Section IV.2 below. On the theoretical side, there have been advances in ab initio calculations, where nucleons are the fundamental degrees of freedom interacting among themselves and with the external electromagnetic probe, see for example [2023FrP....1129094E]. Finally, Bayesian analysis methods have been used to model the energy dependence of the S-factor, starting from the ab initio predictions and applying a polynomial approximation to it.
We review these updates and provide recommended values and uncertainties of the S-factor over the energy range of interest for Solar fusion, based upon a Bayesian averaging of various models.
@@ IV.2 Data sets used for the present review
The 2H(,)3He reaction has a Q value of 5.5 MeV and proceeds through the radiative capture of a proton on deuterium. Different experimental approaches were followed to measure its cross section. 2019EPJA...55..137T irradiated deuterated titanium targets with a proton beam and detected the -rays with two high-purity germanium (HPGe) detectors placed at different angles. The final S-factor is provided at four energies in the 47 - 210 keV range, with approximately 15% statistical uncertainty. More recently, the LUNA Collaboration performed a measurement in the energy range of keV at the underground in the Gran Sasso Laboratories, exploiting the six orders of magnitude suppression of the cosmic radiation background [2021FrASS...7..119Z, 2018IJMPA..3343010C]. A windowless deuterium gas target was used and the -rays emitted by the 2H(,)3He reaction were detected by a large HPGe detector at 90∘ with respect to the beam axis. Great care was taken to minimize all sources of systematic uncertainties in the S-factor overall uncertainty at the 3% level [2020EPJA...56..144M, 2020Natur.587..210M]. These new results provided stringent constraints on cosmological parameters obtained by comparing the precise primordial deuterium abundance predictions of the standard BBN model with astronomical observations [2018ApJ...855..102C]. A deeper discussion of the LUNA results and their implications can be found in [2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y, 2021ApJ...923...49M] 666After the closing of the present review, the 2020Natur.587..210M data were re-analyzed to extract the experimental 2H()3He -ray angular distribution [2024PhRvC.110c2801S]. The data confirm the ab initio angular distributions used in the original analysis by 2020Natur.587..210M.
Finally, a new measurement was performed at the Helmholtz-Zentrum Dresden-Rossendorf in the 300-1000 keV energy range using implanted deuterium targets on tantalum backings and two HPGe detectors [2021PhRvC.103d5805T]. The resulting S-factors show tension with the analysis of the LUNA results extrapolated to keV [2020Natur.587..210M]. However, they are affected by large systematic uncertainties. A new measurement of the reaction is planned at the Felsenkeller laboratory in Germany [2019sone.conf..249B].
@@
In addition to the new accelerator-based results, two recent sets of measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] have also been performed using the inertial confinement fusion plasma-based platform [2017PhPl...24d1407G, 2022FrP....10.4339M, 2023FrP....1180821G], which has recently begun to be exploited for this type of work [e.g., 2016PhRvL.117c5002Z, 2017NatPh..13.1227C]. Both measurements, which were performed at the OMEGA laser facility (see section LABEL:subsec:WG8:Plasma), used laser-driven implosions of spherical plastic-shell capsules filled with H2D2 gas, and measured the emitted -rays using a gas-Cherenkov-detector [2014RScI...85kE124H] that was calibrated applying the technique described in 2019RScI...90l3504Z. The initial experiment [2020PhRvC.101d2802Z] obtained good statistics at an energy of 16 keV by making several repeated measurements, with a final statistical uncertainty of 6% and a systematic one of 17% (dominated by uncertainty in the absolute calibration of the detector). The second experiment [2022FrP....10.4339M] obtained data at three different energies in the 17-37 keV region, with comparable systematic uncertainty, but with larger statistical uncertainty due to fewer repeated measurements. The results obtained on this unique platform agree within error bars with the accelerator-based measurements.